# Solving Truly Massive Budgeted Monotonic POMDPs with Oracle-Guided Meta-Reinforcement Learning

### v1

@@ 4.1 Oracle-Guided RL for Budget-Constrained Monotonic POMDPs
The oracle policy is denoted as $\pi_{\text{oracle}}$ and is obtained by solving the corresponding MDP using value iteration. For a single-component monotonic POMDP with budget $B$, the corresponding MDP has an action space $\mathcal{A}_{\text{MDP}}=\{d,m\}$, identical transition probabilities as the POMDP, and full observability of the state. We use the Proximal Policy Optimization (PPO) algorithm (Schulman et al. 2017) in conjunction with the oracle policy to solve the single-component budget-constrained POMDP. Doing so allows us to treat the problem as that of determining the optimal inspection policy. At each time step, the PPO agent decides whether to perform inspection or not. If the agent decides to take the inspection action, it receives accurate information about the true state of the POMDP. On the other hand, if the agent decides not to inspect, then the action is chosen according to the oracle policy. Hence, the action space for the PPO agent is reduced to $\{q,\neg q\}$. Since the full state is not observable in a POMDP, we utilize the belief $b_{s}$ for planning. The agent’s belief of the true state is updated at each time step using a particle filter approach. For our work, we empirically observe that using the expected belief $\bar{b_{s}}$ and the variance of the belief $\sigma^{2}_{b_{s}}$ suffices for planning.
@@
Hence, for the proposed oracle policy-guided PPO agent, the state at time instant $k$ is given by the vector $[\bar{b}_{s_{k}},c_{k},\sigma^{2}_{b_{s_{k}}}]$. Furthermore, the reward function is defined as follows:
@@
[…]
4.2 Random Forest Approach for Optimal Budget Allocation
In the previous subsection, we detailed the methodology for deriving the optimal policy for a single-component budget-constrained monotonic POMDP. We now extend this to an $n$-component POMDP described in Section 3. Each component in this multi-component POMDP operates independently in terms of transition probabilities, but they share a common total budget, rendering them weakly coupled. While reinforcement learning algorithms have made significant advances, they often face challenges when scaling to the extremely large state and action spaces characteristic of multi-component systems (Sutton and Barto 2018). In the case of a budget-constrained $n$-component POMDP, directly applying a single-component solution approach would result in an exponentially large state space and action space, making such an approach impractical for large-scale problems. To address this scalability issue, we propose an a priori budget distribution strategy to decouple the components, thereby reducing the size of the state and action spaces. Under certain assumptions, the expected maximal time to reach the absorbing state, $\mathbb{E}[T_{\max}]$, for a budget-constrained single-component monotonic POMDP, is concave with respect to the allocated budget $b$ (Vora et al. 2023). For a given component $i$ and allocated budget $b^{i}$, let $\tilde{T}^{i}_{\max}$ be the approximate expected maximal time as a function of the budget, modeled using an exponential function:
$\tilde{T}^{i}_{\max}=\alpha^{i}e^{\beta^{i}b^{i}}+\gamma^{i},$ (3)
where $\alpha^{i},\beta^{i},\gamma^{i}\in\mathbb{R}$ are constants. While many other concave functions could be used to model $\tilde{T}^{i}_{\max}$, we empirically observe that the exponential function provides a good fit for the data. We use a random forest regressor (Breiman 2001) to estimate the parameters of this exponential function. The training dataset for this model is obtained via non-linear least squares regression on multiple $(\mathbb{E}[T_{\max}],b)$ pairs for various budget-constrained single-component monotonic POMDPs. The input to this model includes specific statistics related to the POMDP’s transition function, which are the expected time to reach state 0 without repairs, $\mathbb{E}[T]$, and the variance of this expected time, $\sigma^{2}_{\mathbb{E}[T]}$, as well as the cost of the various actions.
For an $n$-component POMDP with total budget $B$, let the budget allocated to the $i$-th component be $b^{i}$, and the approximate expected maximal time to reach the absorbing state for this budget be $\tilde{T}^{i}_{\max}$.
We then formulate the following constrained maximization problem:
$\max_{b^{i}}\sum_{i=1}^{n}\tilde{T}^{i}_{\max}$ (4)
[…]
Solving (4) yields the approximately optimal budget allocation among the individual components. Since we assume $\tilde{T}^{i}_{\max}$ to be concave in $b^{i}$, the optimization problem becomes a concave maximization problem and is thus guaranteed to have a computationally tractable solution.
@@ 4.3 Optimal Policy for Multi-Component Monotonic POMDPs
We now integrate the approaches described in the previous subsections to compute the optimal policy for an $n$-component POMDP, where $n$ is substantially large. Utilizing the random forest regressor from Section 4.2, we efficiently approximate $\mathbb{E}[T_{\max}]$ for each component $i$. Additionally, we meta-train the oracle-guided PPO agent by constantly updating the policy network’s parameters over a randomly selectied a subset of components and budget values. This approach allows the agent to generalize across components. This meta-trained agent is then utilized to derive the optimal policy $\pi^{i^{*}}$ for each component $i$, following the optimal budget allocation obtained from (4). Consequently, the overall policy for the multi-component POMDP is:
@@ 5 Implementation and Evaluation
We consider an administrative building comprising 1000 infrastructure components, including roofing elements, water fountains, lighting systems, and boilers. Each component’s health is quantified by the Condition Index (CI) (Grussing, Uzarski, and Marrano 2006), which ranges from 0 to 100. For each infrastructure component, we utilize historical CI data to generate the transition probabilities for the corresponding POMDP, modeled using the Weibull distribution (Grussing, Uzarski, and Marrano 2006). We use the weibull_min class from the scipy.stats module in Python to simulate the CI transitions over time. While a seed can be set using the random_state parameter in weibull_min for reproducibility, we did not set one to preserve the stochastic nature of the CI transitions. The condition index deteriorates stochastically over time, influenced by various factors, and can only be accurately assessed through explicit inspections, which incur a cost. A component is considered to have failed when its CI falls below a failure threshold, which is assumed to be 0. Components can be repaired to increase their CI. The building is allocated a maintenance budget of $B=$ 500,000 units for a given horizon of 100 decision steps. At the beginning of the horizon, the CI of all components is 100. The objective of the agents is to maximize the time until failure of the components by efficiently allocating the budget among the components and performing repairs and inspections as needed. The replacement costs (ranging from 50 to 500 units) and inspection costs (ranging from 1 to 5 units) of these components are derived from industry averages. Consistent with the approach described in Section 4.1, we model this objective as a POMDP (with $\alpha=10^{-3}$ in the reward function). This POMDP has roughly $10^{2000}$ states and $3^{1000}$ actions.
5.1 Analysis of Maintenance Policy
@@
    A heuristic policy often used in practice (Lam and Yeh 1994; Straub 2004) where the agent performs inspections at regular intervals and repairs the component when its expected belief about the Condition Index (CI) falls below a predefined threshold. We chose an inspection interval of 5 steps and a repair threshold of 15 after extensive experiments with intervals ranging from 1 to 10 steps and repair thresholds from 5 to 50.
@@
5.2 Analysis of Budget Allocation
Next, we demonstrate the effectiveness of our random forest-based budget allocation strategy. We compare it with a baseline approach that allocates budgets proportional to the ratio of a component’s replacement cost to its $\mathbb{E}[T]$. For a component $i$, we model its $\mathbb{E}[T^{i}_{\max}]$ using $\tilde{T}^{i}_{\max}$ as given in (3). The parameters $\alpha^{i}$ and $\gamma^{i}$ can be estimated directly by considering the boundary conditions: $\gamma^{i}$ is estimated by substituting $b^{i}=0$, representing the scenario where no budget is available, and $\alpha^{i}$ is determined by substituting $b^{i}=\infty$, corresponding to the scenario of unlimited budget, where the supremum of $T^{i}_{\max}$ ($\sup_{b^{i}}T^{i}_{\max}=H=100$) is reached. We then train a random forest regressor to estimate parameter $\beta^{i}$. The training dataset is created by performing non-linear least squares regression on 11 distinct ($T^{i}_{max},b^{i}$) pairs each for 800 components. These pairs correspond to the run-averaged $T^{i}_{\max}$ values and the respective budget values $b^{i}$ from Section 5.1. The input to the random forest model is a vector consisting of the shape and scale factors of the Weibull distribution, which represent $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, along with the replacement and inspection costs for a given component $i$. If a different distribution were used to model the transition probability, we would similarly extract the parameters, $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, for inclusion in the input vector.
@@ 5.2 Analysis of Budget Allocation
Figure 4 shows the prediction performance of the random forest model for a test dataset of 200 components which were not encountered during training. We see that most points on the plot are very close to the perfect prediction line and bad predictions are few in number (29 out of 200 for error threshold of $10^{-4}$). The random forest model achieves a mean squared error (MSE) = $1.8\times 10^{-8}$ for this test dataset. Note that the non-linear least squares regressor constrains $\beta^{i}$ to be $\leq 0$ and hence for some components we observe that $\beta^{i}=0$. We use this trained random forest model to estimate $\tilde{T}^{i}_{max}$ for all 1000 components. Finally, using these approximated expressions, we solve the constrained maximization problem described in (4) to obtain the appropriate budget allocation for the components.
@@
5.3 Analysis of Time Complexity
Finally, we analyze the time complexity of our proposed approach for varying number of components $N$. As mentioned earlier, our method comprises of four major steps:
@@ 5.3 Analysis of Time Complexity
    Random Forest regression for estimating $\tilde{T}^{i}_{max}$ for each component $i$.
@@
    Budget Allocation among components via constrained optimization.
@@
Table 2 presents the times taken for running each of the four processes, with different number of components. The time complexity experiments were performed in Python on a laptop running MacOS with an M2 chip @3.49GHz CPU and 8GB RAM. The times taken for random forest and budget allocation steps are negligible compared to those for performing value iteration and generating optimal policies through meta-PPO. The value iteration is applied to each component independently and hence scales linearly with the number of components. Similarly, Step 4 involves applying the pre-trained policy to each component separately and thus is also linear in the number of components. Consequently, we expect that the time complexity of our algorithm is linear in the number of components, i.e., $O(n)$. This expectation is confirmed by the log-log plot of computational complexity shown in Figure 6. Our algorithm’s performance is thus significantly faster as compared to existing POMDP solvers which would be exponential in the number of states and thus doubly exponential in the number of components (Silver and Veness 2010), (Pineau et al. 2003). If the problem is approached directly as a single POMDP, it will have a prohibitively vast state space of approximately $10^{2000}$ states. Previous work by Vora et al. (2023) demonstrated that standard methods indeed become computationally intractable after a few components due to this combinatorial explosion.

### v2

@@ 4.3 Oracle-Guided RL for a Budget-Constrained Single Component
The oracle policy is denoted as $\pi_{\text{oracle}}$ and is obtained by solving the corresponding MDP using value iteration. For a single-component monotonic POMDP with budget $B$, the corresponding MDP has an action space $\mathcal{A}_{\text{MDP}}=\{d,m\}$, identical transition probabilities as the POMDP, and full observability of the state. We then train a Proximal Policy optimization (PPO) agent Schulman et al. (2017) that queries this oracle selectively: at each time step it chooses either to inspect ($q$) or to defer ($\neg q$), in which case the action recommended by the oracle is executed. Since the full state is not observable in a POMDP, we utilize the belief $b_{s}$ for planning. The agent’s belief of the true state is updated at each time step using a particle filter approach. For our work, we empirically observe that using the expected belief $\bar{b_{s}}$ and the variance of the belief $\sigma^{2}_{b_{s}}$ suffices for planning.
@@
Hence, for the proposed oracle policy-guided PPO agent, the state at time instant $k$ is given by the vector $\bar{b}_{s_{k}},c_{k},\sigma^{2}_{b_{s_{k}}}]$. Furthermore, the reward function is defined as follows:
@@
where $|r_{1}|>|r_{2}|>|r_{3}|$ for all $k$, $0<\alpha<1$ and $\left\lfloor{.}\right\rfloor$ denotes the floor function. This reward function imposes substantial negative rewards for exceeding the budget $B$ and allowing the state $s_{k}$ to reach 0. Additionally, at each time step, the agent receives a positive reward proportional to the time step for maintaining $s_{k}$ above zero and incurs a penalty proportional to the absolute error between the expected belief and the true state. As a result, the agent gets higher rewards for keeping $s_{k}>0$ for a longer time and is heavily penalized when the expected belief deviates significantly from the true state. It is crucial to note that during training, the agent relies solely on the observed reward signals, without access to the true state.
@@ 4.4 Optimal Policy for Multi-Component Monotonic POMDPs
We now integrate the approaches described in Section 4.2 and Section 4.3 to compute the optimal policy for an $n$-component POMDP, where $n$ is substantially large. Utilizing the random forest regressor, we efficiently approximate $\mathbb{E}[T_{\max}]$ for each component ii. Additionally, we meta-train the oracle-guided PPO agent by continuously updating the policy network’s parameters over a randomly selected subset of components and budget values. This approach allows the agent to generalize across components. This meta-trained agent is then utilized to derive the optimal policy $\pi^{i^{*}}$ for each component $i$, following the optimal budget allocation obtained from (6). Consequently, the overall policy for the multi-component POMDP is:
@@ 5 Implementation and Evaluation
In this section, we empirically validate our proposed framework on two disparate domains. The first domain, which we call the infrastructure scenario, involves preventive maintenance for a large-scale building comprising 1000 independent components whose latent condition stochastically degrades over time; our goal is to allocate a finite maintenance budget to maximize the expected survival time of all components. The second domain, the financial loss-budget scenario, addresses portfolio risk management using daily price data for S&P 500 constituents, where each asset is endowed with a debit-only loss budget that depletes under negative returns and can be replenished only through costly recapitalization. In the infrastructure scenario, we compare our oracle-guided meta-PPO approach against baseline heuristics, vanilla PPO, and an idealized oracle policy, reporting results on survival time, cost efficiency, and computational scalability across a range of budget levels. In the financial loss-budget scenario, we focus on analyzing the learned recapitalization policy and assessing the generalizability and window robustness of proposed oracle-guided meta-PPO.
5.1 Implementation and Evaluation for Infrastructure Scenario
@@
We consider an administrative building comprising 1000 infrastructure components, including roofing elements, water fountains, lighting systems, and boilers. Each component’s health is quantified by the Condition Index (CI) Grussing et al. (2006), which ranges from 0 to 100. For each infrastructure component, we utilize historical CI data to generate the transition probabilities for the corresponding POMDP, modeled using the Weibull distribution Grussing et al. (2006). We use the weibull_min class from the scipy.stats module in Python to simulate the CI transitions over time. While a seed can be set using the random_state parameter in weibull_min for reproducibility, we did not set one to preserve the stochastic nature of the CI transitions. The condition index deteriorates stochastically over time, influenced by various factors, and can only be accurately assessed through explicit inspections, which incur a cost. A component is considered to have failed when its CI falls below a failure threshold, which is assumed to be 0. Components can be repaired to increase their CI. The building is allocated a maintenance budget of $B=$ 500,000 units for a given horizon of 100 decision steps. At the beginning of the horizon, the CI of all components is 100. The objective of the agents is to maximize the time until failure of the components by efficiently allocating the budget among the components and performing repairs and inspections as needed. The replacement costs (ranging from 50 to 500 units) and inspection costs (ranging from 1 to 5 units) of these components are derived from industry averages. Consistent with the approach described in Section 4.3, we model this objective as a POMDP (with $\alpha=10^{-3}$ in the reward function). This POMDP has roughly $10^{2000}$ states and $3^{1000}$ actions.
5.1.1 Analysis of Maintenance Policy
@@
    A heuristic policy often used in practice Lam & Yeh (1994); Straub (2004) where the agent performs inspections at regular intervals and repairs the component when its expected belief about the Condition Index (CI) falls below a predefined threshold. We chose an inspection interval of 5 steps and a repair threshold of 15 after extensive experiments with intervals ranging from 1 to 10 steps and repair thresholds from 5 to 50.
@@
5.1.2 Analysis of Budget Allocation
Next, we demonstrate the effectiveness of our random forest-based budget allocation strategy. We compare it with a baseline approach that allocates budgets proportional to the ratio of a component’s replacement cost to its $\mathbb{E}[T]$. For a component $i$, we model its $\mathbb{E}[T^{i}_{\max}]$ using $\tilde{T}^{i}_{\max}$ as given in (5) (see Appendix A for justification of this exponential form). The parameters $\alpha^{i}$ and $\gamma^{i}$ can be estimated directly by considering the boundary conditions: $\gamma^{i}$ is estimated by substituting $b^{i}=0$, representing the scenario where no budget is available, and $\alpha^{i}$ is determined by substituting $b^{i}=\infty$, corresponding to the scenario of unlimited budget, where the supremum of $T^{i}_{\max}$ ($\sup_{b^{i}}T^{i}_{\max}=H=100$) is reached. We then train a random forest regressor to estimate parameter $\beta^{i}$. The training dataset is created by performing non-linear least squares regression on 11 distinct ($T^{i}_{max},b^{i}$) pairs each for 800 components. These pairs correspond to the run-averaged $T^{i}_{\max}$ values and the respective budget values $b^{i}$ from Section 5.1.1. The input to the random forest model is a vector consisting of the shape and scale factors of the Weibull distribution, which represent $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, along with the replacement and inspection costs for a given component ii. If a different distribution was used to model the transition probability, we would similarly extract the parameters, $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, for inclusion in the input vector. Figure 4 shows the prediction performance of the random forest model for a test dataset of 200 components which were not encountered during training. We see that most points on the plot are very close to the perfect prediction line and bad predictions are few in number (29 out of 200 for error threshold of $10^{-4}$). The random forest model achieves a mean squared error (MSE) = $1.8\times 10^{-8}$ for this test dataset. Note that the non-linear least squares regressor constrains $\beta^{i}$ to be $\leq 0$ and hence for some components we observe that $\beta^{i}=0$. We use this trained random forest model to estimate $\tilde{T}^{i}_{max}$ for all 1000 components. Finally, using these approximated expressions, we solve the constrained maximization problem described in (6) to obtain the appropriate budget allocation for the components.
@@ 5.1.2 Analysis of Budget Allocation
5.1.3 Analysis of Time Complexity
Finally, we analyze the time complexity of our proposed approach for varying number of components NN. As mentioned earlier, our method comprises of four major steps:
  1. 1.
    Random Forest regression for estimating $\tilde{T}^{i}_{max}$ for each component ii.
  2. 2.
    Budget Allocation among components via constrained optimization.
  3. 3.
  4. 4.
@@ 5.1.3 Analysis of Time Complexity
Table 2 presents the times taken for running each of the four processes, with different number of components. The time complexity experiments were performed in Python on a laptop running MacOS with an M2 chip @3.49GHz CPU and 8GB RAM. The times taken for random forest and budget allocation steps are negligible compared to those for performing value iteration and generating optimal policies through meta-PPO. The value iteration is applied to each component independently and hence scales linearly with the number of components. Similarly, Step 4 involves applying the pre-trained policy to each component separately and thus is also linear in the number of components. Consequently, we expect that the time complexity of our algorithm is linear in the number of components, i.e., $O(n)$. This expectation is confirmed by the log-log plot of computational complexity shown in Figure 6. Our algorithm’s performance is thus significantly faster as compared to existing POMDP solvers which would be exponential in the number of states and thus doubly exponential in the number of components Silver & Veness (2010), Pineau et al. (2003). If the problem is approached directly as a single POMDP, it will have a prohibitively vast state space of approximately $10^{2000}$ states. Previous work by Vora et al. (2023) demonstrated that standard methods indeed become computationally intractable after a few components due to this combinatorial explosion.
5.2 Implementation and Evaluation for Financial Loss-Budget Management Scenario
Our second experimental scenario addresses a portfolio risk management task. We use daily price data for the S&P 500 constituent stocks over a two-year window, reserving the final $T=120$ trading days for evaluation and using earlier data for training. The core component of the POMDP is an unobserved latent state defined per component as a debit-only loss budget (health), $s_{t}\in[0,100]$. Each component receives a small loss budget: on a day with a negative return, $s_{t}$ is debited proportionally and decreases; on a non-negative day, $s_{t}$ is unchanged; the state does not self-recover. Health increases only when the agent executes recapitalize. All actions draw from one shared, limited budget, and actions are taken when drift relative to the per-component no-loss floor becomes meaningful. This design is practice-inspired for two reasons. First, because we manage a large number of components, governance and our own policy favor a conservative stance: we avoid repeatedly allocating budget to components with recent serial losses, so the health is debit-only and does not auto-replenish. Second, it follows the risk-budgeting workflow described in Benham & Bebee (2024)—set a budget ex ante, allocate and monitor against a benchmark, and treat material drift as a trigger for action. To make the benchmark operational, we instantiate a per-component no-loss floor: losses are deviations that consume the per-component budget; gains are consistent with the floor and do not raise limits by themselves; replenishment occurs only through recapitalize. For training and evaluation, components are assumed independent.
Actions and Costs: The agent’s action space $\mathcal{A}=\{\text{defer, inspect, recapitalize}\}$ manages the per-component loss budget (health). All actions draw from a shared, limited budget $B$ and follow a strict cost hierarchy $c_{\mathrm{recapitalize}}>c_{\mathrm{inspect}}>c_{\mathrm{defer}}$:
  • Defer: Continue with the current position. Incurs a low, continuous cost $c_{\mathrm{defer}}$ each step. Health remains subject to depletion by negative returns.
  • Inspect: Pay $c_{\mathrm{inspect}}$ to obtain a precise observation of the hidden health $s_{t}$ for the selected component.
  • Recapitalize: Pay the high cost $c_{\mathrm{recapitalize}}$ to rebuild health by resetting $s_{t}$ to $100$. This is the only action that increases health.
Objective and Failure Condition. The agent’s objective is to learn a policy $\pi$ that maximizes its survival time. An absorbing failure state is triggered immediately if any component’s health is exhausted, i.e., $s_{t}\leq 0$. For each day the agent survives, it receives a reward of +1. This setup forces the agent to learn a sophisticated policy that balances the continuous drain from defer costs and market losses against the high, discrete costs of inspection and recapitalization, in order to prolong its survival.
5.2.1 Analysis of Recapitalization Policy
We evaluate our approach on a stock-level loss-budget management task constructed from the S&P 500 universe. Starting from $500$ constituents, we retain the subset with at least $80\%$ daily-price coverage over the preceding three years, yielding $471$ components. As in the infrastructure experiment, we reserve the final $T=120$ trading days for evaluation and use earlier data for model training.
State, actions, and costs.
Each component $j$ is modeled as a single-component monotonic POMDP with an unobserved, debit-only loss-budget (health) $s_{t}^{j}\in[0,100]$. Negative returns debit $s_{t}^{j}$ proportionally; non-negative returns leave $s_{t}^{j}$ unchanged; the state does not self-recover. The action set is $\mathcal{A}=\{\text{defer},\text{inspect},\text{recapitalize}\}$ with a strict cost hierarchy $c_{\mathrm{recapitalize}}>c_{\mathrm{inspect}}>c_{\mathrm{defer}}$. A global budget $B_{\text{tot}}$ is shared across all components. Table  3 presents the values of the various parameters used for the experiments.
Parameter Value
Total budget $B_{\text{tot}}$ 15,00015{,}000
Recapitalization cost $c_{\textsc{recap}}$ $10.0$
Inspection cost $c_{\textsc{insp}}$ $0.5$
Defer cost $c_{\textsc{def}}$ $0.2$
Number of components $471$
Evaluation horizon $T$ $120$ days
Table 3: Cost and budget settings for the stock-level scenario.
Policies compared.
We compare four policies:
@@ 5.2.1 Analysis of Recapitalization Policy
    Oracle: full observability of the health $s_{t}$; recapitalize whenever $s_{t}<20$ (no inspection cost).
@@
    Oracle-guided meta-PPO: the agent chooses inspect vs. defer; upon defer, it executes the oracle’s suggested restorative/default control; upon inspect, it pays $c_{\mathrm{inspect}}$ to reduce belief uncertainty. The agent learns when to buy observations and when to accept uncertainty.
@@
    Baseline (Heuristic): fixed inspect every 55 trading days; if the observed $s_{t}<20$, take recapitalize; otherwise defer.
@@
    Vanilla meta-PPO: trained on the same component–budget pairs as the oracle-guided agent but without oracle shaping.
Budget allocation.
We allocate $B_{\text{tot}}$ across the $471$ components using the same random-forest surrogate procedure as in the maintenance experiment: for each component $i$ we fit a concave surrogate for the map $B\mapsto\mathbb{E}[T^{i}_{\max}(B)]$ and solve a tractable concave maximization to obtain per-component budgets.
Training details.
Both oracle-guided meta-PPO and vanilla meta-PPO are trained for $2\times 10^{6}$ timesteps with Adam step size $10^{-4}$, minibatch size $128$, PPO horizon $T_{\mathrm{PPO}}=2048$, and discount factor $0.95$. For each component and policy we run $100$ simulations and report the component-level average $T^{i}_{\max}$; we then average across all $471$ components to obtain $\hat{T}_{\max}$.
Figure 7: S&P 500 stock-level scenario: average survival time $\hat{T}_{\max}$ under a shared budget $B_{\mathrm{tot}}=15{,}000$ across 471 components. Observed ordering: Oracle >> Oracle-guided meta-PPO >> Baseline >> Vanilla meta-PPO.
Findings.
Under the shared budget $B_{\text{tot}}=15{,}000$ across $471$ components and a $120$-day evaluation horizon, we observe a consistent ordering (see Figure 7): $\textbf{Oracle}>\textbf{Oracle-guided meta-PPO}>\textbf{Baseline}>\textbf{Vanilla meta-PPO}$. The Vanilla meta-PPO tends to conserve budget and rarely recapitalizes, yielding the lowest survival time. The Baseline performs periodic inspections (every 55 trading days) and recapitalizes below the threshold but spends budget indiscriminately and misses urgent cases. By contrast, the Oracle-guided meta-PPO learns when to inspect versus defer and when to act, allocating budget to higher-value opportunities; it reliably outperforms the Baseline and closes a substantial portion of the gap to the Oracle upper bound.
5.2.2 Generalizability and Window Robustness of the Oracle-guided Meta-PPO
Design.
We vary the number of components $N\in\{5,10,20,100,471\}$. For each NN, the policy is trained on rolling $120$-day train windows and evaluated both in-sample (train) and on a held-out test window. We report average survival time (days) over r=5r{=}5 seeds; error bars denote $\pm 1$ standard deviation across seeds.
Figure 8: Oracle-guided meta-PPO: train vs. test across component set size NN. The $y$-axis is average survival time (days); the $x$-axis is the number of components.
Findings.
(1) Both train and test curves decrease as NN grows, reflecting budget dilution and increased problem complexity. (2) Train performance is consistently above test with a modest generalization gap that tends to widen at larger NN. (3) Variability is non-negligible and generally larger at higher NN.
Takeaway.
As can be seen from Figure 8, the oracle-guided meta-PPO exhibits window robustness: trends are consistent across train windows, and the train-to-test drop remains moderate. At small NN, the effective exploration/interaction budget is limited, which can hinder learning; as NN increases, richer allocation opportunities make better use of the oracle guidance even though absolute survival time declines under a fixed total budget.